OpenStudy (anonymous):

y=x^1/3 on what interval is this function continuous? the answer is (-) infinity to (+) infinity. however how can you take a cubed root of 0?

OpenStudy (anonymous):

no cube for 0

OpenStudy (anonymous):

You can, since: \[ 0^3=0\cdot0\cdot0=0 \]So, it is continuous for: \[ x\in\mathbb{R} \]

OpenStudy (anonymous):

thats what i was assuming but with a power of 1/3 that insinuates cubed root. and you cant take root of 0. does that rule not apply if it is written as 1/3

OpenStudy (anonymous):

Nope, sorry, here's the definition of a cube root: \[ \sqrt[3]{a}=b\\ a\in\mathbb{R} \]Is the solution to the equation: \[ b^3=a\\ b\in\mathbb{R} \](That's the typical definition, you could also show using roots of unity that the other two solutions are the case, but these are complex)

OpenStudy (anonymous):

So, 0 is the solution to the equation: \[ \sqrt[3]{0}=a\implies a^3=0 \]But we know: \[ 0^3=0 \]And \[ 0\in\mathbb{R} \]So: \[ a=0=\sqrt[3]{0} \]